Find the mass of a triangular lamina whose vertices are (0,0), (1, 1), (0.1). if the density of the lamina is p(x, y) = sin (y^2).


1
Expert's answer
2021-02-24T07:59:19-0500

Consider the triangular lamina as shown in the figure below:






The region "R" in cartesian coordinates is given by,


"R=\{(x,y)|0\eqslantless y \eqslantless1, 0 \eqslantless x \eqslantless y \}"


The mass of the lamina with density "\rho(x,y)=sin(y^2)" is evaluated as,


Mass"(m)=\iint_{R}\rho(x,y)dA"


"=\int_{y=0}^{1}\int_{x=0}^{y}sin(y^2)dxdy"


"=\int_{y=0}^{1}sin(y^2)[x]_{x=0}^{y}dy"


"=\int_{y=0}^{1}ysin(y^2)dy"


Let "u=y^2" , then "\frac{1}{2}du=ydy" , "0\eqslantless u \eqslantless 1" , so the integral is,


"=\int_{u=0}^{1}\frac{1}{2}sin(u)du"


"=\frac{1}{2}[-cos(u)]_{=0}^{1}"


"=\frac{1}{2}(1-cos(1)) \approx 0.22984"


Therefore, the mass of the lamina is "m=\frac{1}{2}(1-cos(1)) \approx 0.22984" .